Circular No. 8642

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2005 X1 (BESHORE) Edward C. Beshore has discovered a diffuse comet with slight central

condensation (diameter ∼ 15′′; possible broad, faint tail ∼ 5′′ long in p.a. ∼ 300^{o}) on Catalina Sky Survey images (0.68-m Schmidt telescope; discovery observation tabulated below). Following posting on the ‘NEO Confirmation Page’, Š. Gajdoš and J. Vilagi confirm the diffuse nature of the object (coma diameter ∼ 8′′) on CCD images taken on Dec. 9.1 UT with the 0.6-m f/5.5 reflector at Modra; G. Hug (Eskridge, Kansas, 0.7-m reflector) writes that his images taken on Dec. 9.4 also show the object to be diffuse.

2005 UT  α_{2000}  δ_{2000}  Mag.
Dec. 7.41028  10^{h}13^{m}15.^{s}52  +38^{o}35′32′′.9  19.6

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2005-X51.

T = 2005 Sept.30.861 TT  ω = 142.^{o}571
Ω = 301.419 }2000.0
q = 3.27707 AU  i = 96.463

V445 PUPPIS D. K. Lynch, R. J. Rudy, C. C. Venturini, and S. Mazuk, Aerospace

Corporation; R. C. Puetter, University of California at San Diego; R. B. Perry, Langley Research Center, NASA; and B. Walp, Lick Observatory, report 1.0- to 2.5-μm spectroscopy of V445 Pup (cf. IAUC 7559, 7580) with the Lick 3.0-m telescope (+ VNIRIS) at Nov. 16.535 UT. The object has faded and the thermal dust emission (Lynch et al. 2004, A.J. 128, 2962) has virtually disappeared. No visible counterpart has appeared, so it is likely that the dust has cooled significantly. Structured He I emission at 1.0830 μm remains strong.

COMET C/2005 W2 (CHRISTENSEN)

Improved elliptical orbital elements from MPEC 2005-X23:

T = 2006 Mar. 28.0034 TT  ω = 111.^{o}8212
e = 0.825331  Ω = 336.5983 }2000.0
q = 3.330897 AU  i = 11.2651

a = 19.069727 AU n^{o} = 0.0118355 P = 83.3 years

2005 December 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green